Meeta saves ₹ 4000 from her salary. If this is 10% of her salary. What is her salary? Solution:-

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume Meeta’s salary be ₹x,
Then,
10% of₹x =₹4000
(10/100) × (x) = 4000
X = 4000 × (100/10)
X = 4000 × 10
X =₹40000
∴Meeta’s salaryis₹40000.

Thanks